# Maintenance Optimization for Asset Networks with Unknown Degradation Parameters

### v2

@@ 6.3 Numerical results and model validation
In this section, we assess the predictive accuracy and robustness of the best solution approaches identified in Tables 6, 7 and 8 by evaluating their performance directly on the test data. Specifically, we analyze the ability of the model to generalize beyond the training data and evaluate how well the estimated hyperparameters of the distributions $\Lambda$ and $\Phi$ capture the degradation behavior of other X-ray tubes across different case study instances. To this, the time series in the test set $\mathcal{I_{\text{test}}}$ are transformed into BMDP trajectories. All performance results are derived from $10^{4}$ repetitions. The length of each repetition is set to $10^{3}$, with trajectories sampled from $\mathcal{I_{\text{test}}}$ with replacement. The reported half-widths correspond to asymptotic $95\%$ confidence intervals.
@@ 7 Conclusion
In this paper, we developed a novel maintenance model aimed at optimizing the management of a network of advanced industrial assets such as medical imaging devices and wind turbines, which are prone to costly and disruptive unplanned downtimes. By addressing the limitations of traditional condition-based maintenance models, which often assume homogeneous assets in isolation, our approach introduces a more practical framework that accounts for component heterogeneity and economic dependence between assets.
@@
Our approach leverages real-time degradation data to learn cost-effective maintenance strategies for asset networks with economic dependencies using a partially observable Markov decision process (POMDP) framework. We further addressed the computational challenges of solving POMDPs by employing a deep reinforcement learning (DRL) approach, enabling the derivation of near-optimal maintenance policies under parameter uncertainty. The DRL-based approximate policy iteration algorithm demonstrated its effectiveness by learning complex opportunistic maintenance strategies, directly improving upon common heuristic methods.
Through theoretical contributions, including the establishment of the structural properties of optimal replacement policies of the underlying Markov decision process (MDP) model and the reformulation of the POMDP as a Bayesian MDP (BMDP), our approach facilitates scalable and efficient maintenance solutions for industrial-scale asset networks. The practical value of the proposed model was highlighted through a case study on degrading interventional X-ray system components, providing actionable managerial insights.
Overall, this work contributes to the literature by advancing maintenance strategies for asset networks under uncertainty, demonstrating the potential to significantly reduce unplanned downtime, improve cost efficiency, and enhance operational reliability across industries.
@@ Acknowledgements
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-5192. This work was supported by the Netherlands Organisation for Scientific Research (NWO). Project: NWO Big data - Real Time ICT for Logistics. Number: 628.009.012. The work of Stella Kapodistria is supported by NWO through the Gravitation-grant NETWORKS-024.002.003.
@@ Appendix A Theorems and proofs
We provide the proofs of the established properties of the optimal replacement policy of the underlying MDP model presented in Section 3.5, inspired by the proof of Soltani et al. (2024, Theorem 3.7).
@@
Firstly, note that $\mathcal{U}(y)\subseteq\mathcal{U}(x)$ and that $a^{*}(x)\in\mathcal{U}(y)$. For $a\in\mathcal{U}(x)$, define the following sets:
@@
$C(y,a)-C(y,a^{*}(x))$ $=c^{\textrm{ST}}\left(\mathds{I}\{a>0\}-\mathds{I}\{a^{*}(x)>0\}\right)+\sum_{m\in\mathcal{G}(x,a)}c_{m}^{\textrm{PM}}-\sum_{m\in\mathcal{J}(x,a)}c_{m}^{\textrm{PM}}$
@@
Denote with $\hat{y}(a)$ and $\hat{x}(a)$ the next state after taking action $a\in\mathcal{U}(y)$ in state $y$ resp. $x$. Since we do maintenance on all machines with a more severe degradation level, we have that
$\hat{y}(a^{*}(x))\overset{d}{=}\hat{x}(a^{*}(x)).$
@@
for all non-decreasing functions $\phi:\mathbb{R}\rightarrow\mathbb{R}$ for which these expectations exist (Shaked and Shanthikumar, 2007, 1.A.7). Note that it suffices to have this property for one state variable as the argument can be repeated for other state variables. In particular, we can show that this property holds for the value function.
Let $V(\cdot)$ denote the value function of an optimal policy that solves Eq. (1) and let $V^{n}(\cdot)$ denote the value function at the $n$-th iteration of the value iteration algorithm, i.e., the value iteration algorithm produces the sequence $\{V^{n}(\cdot)\}_{n\in\mathbb{Z_{+}}}$. Since our state space is a Borel space, $\gamma\in(0,1)$ and costs are bounded from above, the value iteration algorithm is guaranteed to converge point-wise to the optimal value function of an optimal policy (Bertsekas and Shreve, 1996, Proposition 9.14). That is, for all states $h$ of the underlying MDP, $V^{n}(h)\rightarrow V(h)$ for any arbitrary starting position through the value iteration algorithm.
@@
To avoid notational clutter, we will omit the degradation process parameters $\lambda_{1},\phi_{1},\lambda_{2}$, and $\phi_{2}$ in the following argument. Let $A(t)=(X_{1}(t),X_{2}(t))$ denote the random amount of damage accumulated per machine after $t$ time units (without maintenance intervention). Let $s_{0}=(0,0)$ be an arbitrary healthy state. The Bellman optimality equations for the case $M=2$ read as follows:
@@
All the constants that appear on the RHS and $A(1)$ are non-decreasing in $x_{1}$, therefore, $V^{n+1}((x_{1},x_{2}))$ is non-decreasing in $x_{1}$. All in all, this implies that
@@
We proceed with proving that for all $a\in\mathcal{U}(y)$,
@@
To do this, we expand the LHS using a one-step argument and apply the results from Eq. (10) and Eq. (11).
@@
Let $x=(x_{1},\lambda_{1},\phi_{1},\ldots,x_{M},\lambda_{M},\phi_{M})$ denote a state. Let $\mathcal{M}(x)\subseteq\mathcal{M}$ again the set of machines for which it is optimal to do maintenance in state $x$. For any $x^{\prime}=(x_{1},\lambda^{\prime}_{1},\phi^{\prime}_{1},\ldots,x_{M},\lambda^{\prime}_{M},\phi^{\prime}_{M})\in\mathcal{S}$ that represents a state with worse degradation parameters than $x$, i.e.,
@@
In this case, $\mathcal{U}(x^{\prime})=\mathcal{U}(x)$ and thus $a^{*}(x)\in\mathcal{U}(x^{\prime})$. Since states $x$ and $x^{\prime}$ have identical degradation levels and differ only in their degradation process parameters, the costs of taking any action are the same. Therefore, for all $a\in\mathcal{U}(x)$,
@@
Denote with $\hat{x}^{\prime}(a)$ and $\hat{x}(a)$ the next state after taking action $a\in\mathcal{U}(x)$ in state $x^{\prime}$ resp. $x$. Since we do maintenance on all machines with “worse” degradation process parameters, we have that
$\hat{x}^{\prime}(a^{*}(x))\overset{d}{=}\hat{x}(a^{*}(x)).$
The assumptions $\lambda^{\prime}_{m}\geq\lambda_{m}$ and $Y^{\prime}_{m}\geq_{\textrm{st}}Y_{m}$ for all $m\in\mathcal{M}(x)$ ensure that $\hat{x}^{\prime}(a)\geq_{\textrm{st}}\hat{x}(a)$ for all $a\in\mathcal{U}(x)$ (Shaked and Shanthikumar, 2007, Theorem 1.A.4). A similar argument to the one used in the proof of Theorem 1, which uses induction on the iterations of the value iteration algorithm, shows that the value function is non-decreasing with respect to the degradation process parameters $\lambda_{m}$ and $\phi_{m}$. This means that
@@
We expand again the LHS using a one-step argument and use the previous established results from Eq. (12) and Eq. (13).
@@
$a^{*}(x)\in\textrm{arg}\underset{a\in\mathcal{A}(x^{\prime})}{\textrm{min}}Q(x^{\prime},a).$
@@
$a^{*}(x^{\prime})\in\textrm{arg}\underset{a\in\mathcal{A}(x^{\prime})}{\textrm{min}}Q(x^{\prime},a).$
Meaning that
@@ Appendix B Nomenclature
Appendix C Deep reinforcement learning hyperparameters
We present the selected hyperparameters for the training algorithm below. All neural networks employ the rectified linear unit (ReLU) activation function.

### v3

@@ 6.3 Numerical results and model validation
In this section, we assess the predictive accuracy and robustness of the best solution approaches identified in Tables 6, 7 and 8 by evaluating their performance directly on the test data. Specifically, we analyze the ability of the model to generalize beyond the training data and evaluate how well the estimated hyperparameters of the distributions $\Lambda$ and $\Phi$ capture the degradation behavior of other X-ray tubes across different case study instances. To this end, the time series in the test set $\mathcal{I_{\text{test}}}$ are transformed into BMDP trajectories. All performance results are derived from $10^{4}$ repetitions. The length of each repetition is set to $10^{3}$, with trajectories sampled from $\mathcal{I_{\text{test}}}$ with replacement. The reported half-widths correspond to asymptotic $95\%$ confidence intervals.
@@ 7 Conclusion
In this paper, we developed a novel maintenance model aimed at optimizing the management of a network of advanced industrial assets prone to costly and disruptive unplanned downtimes. By addressing the limitations of traditional condition-based maintenance models, which often assume homogeneous assets in isolation, our approach introduces a more practical framework that accounts for both asset heterogeneity and component heterogeneity, as well as economic dependencies between assets. For tractability purposes, the current formulation does not incorporate resource constraints. However, it can be extended to include shared resources—such as repair crews or spare parts—thereby allowing its application to asset networks with additional logistical complexities. Moreover, the degradation dynamics are modeled as a compound Poisson process with independent and identically distributed shock sizes. While this choice is standard in the reliability literature and supported by our case study data, it abstracts from certain practical complexities. In particular, we assume independent shock processes across assets and stationarity of degradation parameters within a component’s lifetime, thereby excluding, for example, common-cause or environment-driven correlations and non-stationary or regime-switching degradation behavior. Extending the framework to correlated or non-stationary degradation processes constitutes a promising direction for future research.
@@
Using a partially observable Markov decision process (POMDP) framework, our approach leverages real-time degradation data to learn cost-effective maintenance strategies for asset networks with economic dependencies. We further addressed the computational challenges of solving POMDPs by employing a deep reinforcement learning (DRL) approach, enabling the derivation of near-optimal maintenance policies under parameter uncertainty. The DRL-based approximate policy iteration algorithm demonstrated its effectiveness by learning complex opportunistic maintenance strategies, directly improving upon common heuristic methods.
Through theoretical contributions, including the establishment of the structural properties of optimal replacement policies of the underlying Markov decision process (MDP) model and the reformulation of the POMDP as a Bayesian MDP (BMDP), our approach facilitates scalable and efficient maintenance solutions for industrial-scale asset networks. However, the scope of the established structural properties is limited to the full-information setting and does not directly transfer to the partial-information case. Nevertheless, we expect that the monotonicity result in the degradation level also holds under partial information, and a similar monotonicity result could be established for another state variable, namely the operational age, which we leave for future work. Such structural properties could aid the development of heuristic methods tailored to the BMDP case.
The practical value of the proposed model was highlighted through a case study on degrading interventional X-ray system components, providing actionable managerial insights. The instances considered in this work are relatively small and assume asset-level homogeneity (i.e., all component replacements stem from the same pool), and therefore require relatively few samples to train our DRL algorithm; industrial-scale problems require many more, which can make single-node execution potentially prohibitive. By leveraging distributed computing and high-performance clusters, sample collection can be parallelized, keeping computation times manageable even for large, complex industrial-scale instances. Our approach benefits significantly from smart initialization heuristics; developing a scalable initialization method for cases where heterogeneity is present at both the asset and component levels therefore appears promising.
However, when no computationally convenient parametric form exists (i.e., prior-to-posterior belief updating is not tractable) to model component heterogeneity, exact reformulation of the maintenance problem as a BMDP is no longer feasible. Alternative inference techniques, such as variational inference or Monte Carlo sampling methods, introduce approximations and significant computational challenges in the DRL algorithm. Consequently, it is not clear what the effect is on the trained DRL-based policies using such techniques. As a workaround, we proposed a solution where policies trained on the underlying MDP are applied in a POMDP setting using an open-loop feedback approach. However, our numerical experiments indicate that this solution is effective only in controlled settings and not when component heterogeneity is highly volatile, presenting an opportunity for future work.
Finally, the performance results of the trained policies under different information levels provide insight into the value of increasing the information level. In practice, additional information is obtained through investments in data availability, quality and processing infrastructure. The value of information can therefore be interpreted as a trade-off between such investment costs and the potential for improved CBM planning. Asset managers can evaluate this trade-off through scenario analysis to support investment decisions, which is a promising direction for future research.
@@ Acknowledgements
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-5192. This work was supported by the Netherlands Organisation for Scientific Research (NWO). Project: NWO Big data - Real Time ICT for Logistics. Number: 628.009.012. The work of Stella Kapodistria is supported by NWO through the Gravitation-grant NETWORKS-024.002.003. Collin Drent received support from NWO through Grant VI.Veni.241E.058.
@@ Appendix A Theorems and proofs
We provide the proofs of the established properties of the optimal replacement policy of the underlying MDP model presented in Section 3.5, inspired by the proof of Soltani et al. [2024, Theorem 3.7]. Note that we restrict attention to the case of $M=2$ parallel assets; however, the arguments extend analogously to the case $M>2$.
@@
Firstly, note that $\mathcal{U}(y)\subseteq\mathcal{U}(x)$ and that $a^{*}(x)\in\mathcal{U}(y)$. To prove the statement, we use a one-step argument to expand the LHS and show that, for all $a\in\mathcal{U}(y)$, the RHS is nonnegative:
$Q(y,a)-Q(y,a^{*}(x))$ $=\underbrace{C(y,a)-C(y,a^{*}(x))}_{\text{cost difference}}+\underbrace{\gamma\mathbb{E}[V(\hat{y}(a))]-\gamma\mathbb{E}[V(\hat{y}(a^{*}(x)))]}_{\text{future-value difference}}.$
We proceed with deriving bounds for each bracketed term on the RHS.
For $a\in\mathcal{U}(x)$, define the following sets:
@@
$C(y,a)-C(y,a^{*}(x))$ $=c^{\textrm{ST}}\left(\mathds{1}_{\{\sum_{m\in\mathcal{M}}a_{m}>0\}}-\mathds{1}_{\{\sum_{m\in\mathcal{M}}a_{m}^{*}(x)>0\}}\right)+\sum_{m\in\mathcal{G}(x,a)}c_{m}^{\textrm{PM}}-\sum_{m\in\mathcal{J}(x,a)}c_{m}^{\textrm{PM}}$
@@
This implies Eq. 10.
Denote with $\hat{y}(a)$ and $\hat{x}(a)$ the next state after taking action $a\in\mathcal{U}(y)$ in states $y$ and $x$, respectively. Since we do maintenance on all machines with a more severe degradation level, we have that
$\hat{y}(a^{*}(x))\overset{\textrm{d}}{=}\hat{x}(a^{*}(x)).$
@@
for all non-decreasing functions $\phi:\mathbb{R}\rightarrow\mathbb{R}$ for which these expectations exist [Shaked and Shanthikumar, 2007, 1.A.7]. Note that it suffices to have this property for one state variable as the argument can be repeated for other state variables. In particular, we can show that this property holds for the value function.
Let $V(\cdot)$ denote the value function of an optimal policy that solves Eq. (1) and let $V^{n}(\cdot)$ denote the value function at the $n$-th iteration of the value iteration algorithm, i.e., the value iteration algorithm produces the sequence $\{V^{n}(\cdot)\}_{n\in\mathbb{Z}_{+}}$. Since our state space is a Borel space, $\gamma\in(0,1)$ and costs are bounded from above, the value iteration algorithm is guaranteed to converge point-wise to the optimal value function of an optimal policy [Bertsekas and Shreve, 1996, Proposition 9.14]. That is, for all states $h$ of the underlying MDP, $V^{n}(h)\rightarrow V(h)$ for any arbitrary starting position through the value iteration algorithm.
@@
To avoid notational clutter, we will omit the degradation process parameters $\lambda_{1},\phi_{1},\lambda_{2}$, and $\phi_{2}$ in the following argument. Let $D(t)=(X_{1}(t),X_{2}(t))$ denote the random amount of damage accumulated per machine after $t$ time units (without maintenance intervention). Let $s_{0}=(0,0)$ be an arbitrary healthy state. The Bellman optimality equations for the case $M=2$ read as follows:
@@
All the constants that appear on the RHS and $D(1)$ are non-decreasing in $x_{1}$; therefore, $V^{n+1}((x_{1},x_{2}))$ is non-decreasing in $x_{1}$. All in all, this implies that
@@
We proceed by proving that for all $a\in\mathcal{U}(y)$,
@@
This result follows directly from the previously established results in Eq. (10) and Eq. (11):
@@
Let $x=(x_{1},\lambda_{1},\phi_{1},\ldots,x_{M},\lambda_{M},\phi_{M})$ denote a state. Let $\mathcal{M}(x)\subseteq\mathcal{M}$ denote the set of machines for which it is optimal to do maintenance in state $x$. For any $x^{\prime}=(x_{1},\lambda^{\prime}_{1},\phi^{\prime}_{1},\ldots,x_{M},\lambda^{\prime}_{M},\phi^{\prime}_{M})$ that represents a state with worse degradation parameters than $x$, i.e.,
@@
Note that, in this case, $\mathcal{U}(x^{\prime})=\mathcal{U}(x)$ and thus $a^{*}(x)\in\mathcal{U}(x^{\prime})$. The proof follows the same approach as the proof of Theorem 1. Using a one-step argument, we expand the LHS and show that, for all $a\in\mathcal{U}(x^{\prime})$, the RHS is nonnegative:
$Q(x^{\prime},a)-Q(x^{\prime},a^{*}(x))$ $=\underbrace{C(x^{\prime},a)-C(x^{\prime},a^{*}(x))}_{\text{cost difference}}+\underbrace{\gamma\mathbb{E}[V(\hat{x}^{\prime}(a))]-\gamma\mathbb{E}[V(\hat{x}^{\prime}(a^{*}(x)))]}_{\text{future-value difference}}.$
We proceed with deriving bounds for each bracketed term on the RHS.
Since states $x$ and $x^{\prime}$ have identical degradation levels and differ only in their degradation process parameters, the costs of taking any action are the same. Therefore, for all $a\in\mathcal{U}(x)$,
@@
Denote with $\hat{x}^{\prime}(a)$ and $\hat{x}(a)$ the next state after taking action $a\in\mathcal{U}(x)$ in states $x^{\prime}$ and $x$, respectively. Since we do maintenance on all machines with “worse” degradation process parameters, we have that
$\hat{x}^{\prime}(a^{*}(x))\overset{\textrm{d}}{=}\hat{x}(a^{*}(x)).$
The assumptions $\lambda^{\prime}_{m}\geq\lambda_{m}$ and $Y^{\prime}_{m}\geq_{\textrm{st}}Y_{m}$ for all $m\in\mathcal{M}(x)$ ensure that $\hat{x}^{\prime}(a)\geq_{\textrm{st}}\hat{x}(a)$ for all $a\in\mathcal{U}(x)$ [Shaked and Shanthikumar, 2007, Theorem 1.A.4]. A similar argument to the one used in the proof of Theorem 1, which uses induction on the iterations of the value iteration algorithm, shows that the value function is non-decreasing with respect to the degradation process parameters $\lambda_{m}$ and $\phi_{m}$. This means that
@@
This result follows directly from the previously established results in Eq. (12) and Eq. (13):
@@
$a^{*}(x^{\prime})\in\textrm{arg}\underset{a\in\mathcal{U}(x^{\prime})}{\textrm{min}}Q(x^{\prime},a).$
This means that
@@ Appendix B Nomenclature
Nomenclature
  • $\mathcal{M}$
    Set of machines
  • $M$
    Number of machines
  • $m$
    Machine index
  • $t$
    Time step
  • $N_{m}(t)$
    Shock arrival process of machine $m$
  • $\lambda_{m}$
    Shock arrival rate of machine $m$
  • $\Lambda_{m}$
    Sampling distribution of the shock
    arrival rate of machine $m$
  • $\alpha_{m}$, $\beta_{m}$
    Shape and scale hyperparameters of the Gamma prior for $\lambda_{m}$
  • $Y^{(i)}_{m}$
    Shock size of machine $m$
  • $\phi_{m}$
    Shock size distribution parameter of machine $m$
  • $p_{m}$
    Geometric distribution parameter of machine $m$
  • $\Phi_{m}$
    Sampling distribution of the shock size distribution parameter of machine $m$
  • $r_{m}$, $s_{m}$
    Hyperparameters of the conjugate prior for $\phi_{m}$
  • $X_{m}(t)$
    Degradation process of machine $m$
  • $x_{m}(t)$
    Degradation level of machine $m$
  • $\xi_{m}$
    Failure threshold of machine $m$
  • $k_{m}(t)$
    Shocks sustained by machine $m$ in
    period (t1,t](t-1,t]
  • $z_{m}(t)$
    Accumulated damage of machine $m$ in period (t1,t](t-1,t]
  • $t_{m}(t)$
    Machine age of machine $m$
  • $h$
    Underlying MDP state
  • $o$
    Observed state
  • $\tilde{h}$
    BMDP state
  • $\mathcal{U}(h)$
    State-dependent action set
  • $c^{\textrm{PM}}_{m}$
    PM cost of machine $m$
  • $c^{\textrm{CM}}_{m}$
    CM cost of machine $m$
  • $c^{\textrm{ST}}$
    Setup cost
  • $\gamma$
    Discount factor
  • $C(h,a)$
    Cost function
  • $\textbf{L}_{i}$
    Information level $i$
  • $\pi^{\textbf{L}_{i}}$
    Policy under information level $\textbf{L}_{i}$
  • $J(\pi^{\textbf{L}_{i}})$
    Total expected discounted cost under policy $\pi^{\textbf{L}_{i}}$
  • $\tau_{m}^{\textrm{PM}}$
    PM threshold of machine $m$
  • $\tau_{m}^{\textrm{OPM}}$
    OPM threshold of machine $m$
  • $f^{\textbf{L}}(h)$
    Feature vector
Appendix C Examples of exponential-family representations
In this appendix, we present two commonly used discrete shock-size distributions and show how they can be written in canonical exponential-family form with a linear sufficient statistic.
C.1 Binomial distribution
Suppose the shock size is binomially distributed with a fixed number of trials $n_{m}\in\mathbb{N}$ and success probability $p_{m}\in(0,1)$. Its probability mass function is
$f(x\mid p_{m})=\binom{n_{m}}{x}p_{m}^{x}(1-p_{m})^{n_{m}-x},\quad x=0,1,\dots,n_{m}.$
Rewriting the density,
$f(x\mid p_{m})=\binom{n_{m}}{x}\exp\Big(x\ln p_{m}+(n_{m}-x)\ln(1-p_{m})\Big)$
$\binom{n_{m}}{x}\exp\Big(x\ln\!\Big(\frac{p_{m}}{1-p_{m}}\Big)+n_{m}\ln(1-p_{m})\Big).$
Hence, the binomial distribution can be written in canonical exponential-family form
$f(x\mid\phi_{m})=h_{m}(x)\exp\big(\phi_{m}T_{m}(x)-A_{m}(\phi_{m})\big),$
with
$T_{m}(x)=x,\quad\phi_{m}=\ln\!\Big(\frac{p_{m}}{1-p_{m}}\Big),$
$h_{m}(x)=\binom{n_{m}}{x},\quad A_{m}(\phi_{m})=n_{m}\ln\big(1+e^{\phi_{m}}\big).$
Thus, the binomial distribution belongs to the one-parameter exponential family with linear sufficient statistic.
C.2 Poisson distribution
Suppose the shock size is Poisson distributed with mean $\mu_{m}>0$. Its probability mass function is
$f(x\mid\mu_{m})=\frac{e^{-\mu_{m}}\mu_{m}^{x}}{x!},\quad x=0,1,2,\dots.$
Rewriting,
$f(x\mid\mu_{m})=\frac{1}{x!}\exp\big(x\ln\mu_{m}-\mu_{m}\big).$
This can be written in canonical exponential-family form as
$f(x\mid\phi_{m})=h_{m}(x)\exp\big(\phi_{m}T_{m}(x)-A_{m}(\phi_{m})\big),$
with
$T_{m}(x)=x,\quad\phi_{m}=\ln\mu_{m},$
$h_{m}(x)=\frac{1}{x!},\quad A_{m}(\phi_{m})=e^{\phi_{m}}.$
Hence, the Poisson distribution also belongs to the one-parameter exponential family with linear sufficient statistic.
Appendix D Deep controlled learning hyperparameters
We present the selected hyperparameters for the deep controlled learning algorithm below. All neural networks employ the rectified linear unit (ReLU) activation function.